OpenStudy (anonymous):

How do you factor x^2-x-12

OpenStudy (bohotness):

x^2 -x -12 = 0 Separate : ( x^2 +3*x ) + ( -4*x -12 ) = 0 Commutative Law : ( x^2 -4*x ) + ( 3*x -12 ) = 0 Distributive Law : x*( x -4 ) + 3*( x -4 ) = 0 Factor : ( x +3 )*( x -4 ) = 0

OpenStudy (anonymous):

How would you factor it if it had a coefficient in front like 2x^2+5x+6?